# references.bib {refsection}

### v3

@@ 3.2.1 Comparisons of expert and general population survey on AI axes of disparities
Experts were asked to rate how likely different potential axes of disparities would cause AI to perform differently for sub-groups in Africa while Gen Pop. respondents were asked to rate how likely different demographics would cause AI to perform differently for the participant themselves. Results are reported in Fig. 6. Overall most experts (71%) indicated that all demographics were likely or very likely to be axes of AI disparities/AI biases for sub-groups in Africa, with literacy, race, language, rural vs. urban location, and national income level being the top 5 axes of disparities. Most Gen Pop. respondents did not feel that AI tools would perform differently for them compared to other people. In agreement with experts, Gen Pop. indicated that literacy had the highest level of causing an AI tool to perform differently. Given that our respondents were mostly from educated backgrounds, 32% indicated that their literacy level would allow an AI tool for health to perform better for them. Country of residence, country of birth, race, and colonial history of country of origin had the highest percentage of respondents indicating that they would cause an AI tool to perform worse for them (Fig. 6).
@@
Figure 6: Perceptions from a) experts and b) Gen Pop. respondents on how likely different demographics would cause an AI tool to perform differently for people in Africa. Note that Better is not necessarily good, it reflects that our participant population is better off than other people in the same country. e.g., Most of our survey participants are skewed towards high literacy and education levels compared to the general population, and so AI may perform better for them than for others.
@@
THEME: Opportunities for AI to Address Health in Africa
Subtheme Representative quote
THEME: Intersectional Inequities and Systemic Factors Impacting Health Equity in Africa
Subtheme Representative quote
THEME: Social and Ethical Considerations for Developing
and Deploying AI solutions in Health in Africa
Subtheme Representative quote
THEME: Capacity Building for Effective Implementation and Adoption
of AI Solutions in Health in Africa
Subtheme Representative quote
Table 3: A summary of the themes and sub themes with representative quotes from the expert in-depth interviews
3.3 Thematic analysis of expert qualitative interviews
3.3.1 Theme: Trust In AI in Africa
Figure 7: Major theme from the expert IDIs: Trust in AI. Summary of the subthemes illustrating the identified causes for mistrust and approaches for building trust.
Subtheme: Mistrust due to marginalization, colonization, misunderstanding of the technology, and perceived lack of transparency, fairness, and accountability
Mistrust due to marginalization: Experts pointed to skepticism surrounding AI in Africa based on a history of marginalization from the Global North. “a lot of times things that are seen to be coming from the Global North are viewed with a lot of skepticism. -P025 [50-59, Nigeria, Medical Doctor]. This distrust is further fueled by concerns of exploitation by foreign entities and worries of job insecurity among healthcare workers with sentiments, such as “What are the hidden things that they really want that they’re not letting us see?”- P025 [50-59, Nigeria, Medical Doctor], or “people are scared of the power of AI, they think that it is going to wipe their businesses off, is going to take their jobs” - P168 [30-39, Ghana, Educator].
@@ 3.3.1 Theme: Trust In AI in Africa
3.3.2 Theme: Opportunities for AI to Address Health in Africa
@@
Figure 8: Major theme from the expert IDIs: Opportunities for AI to address health in Africa. Summary of the subthemes illustrating the foremost use cases for AI in health and the desired features for AI solutions to improve efficacy and adoption.
@@
3.3.3 Theme: Intersectional Inequities and Systemic Factors Impacting Health Equity in Africa
@@
Figure 9: Major theme from the expert IDIs: Intersectional Inequities and Systemic Factors impacting health equity in Africa. Summary of the subthemes illustrating the existing causes of health inequities and proposed approaches for addressing them.
@@
3.3.4 Theme: Social and Ethical Considerations for Developing and Deploying AI Solutions in Health in Africa
@@
Figure 10: Major theme from the expert IDIs: Social and Ethical considerations for developing and deploying AI solutions in Health in Africa. Summary of the subthemes illustrating the social, ethical, privacy and data-related considerations.
@@
In addition to social considerations, participants emphasized the importance of responsible and ethical implementation of AI solutions to ensure fairness and avoid negative consequences.
Ethical deployment: While the participants acknowledged that AI holds great promise for revolutionizing healthcare in Africa, potentially making it more “accessible, efficient, and equitable” (P115 [21-29, Cameroon, Machine Learning Researcher]), they caution against uncritical adoption, stressing the need for ethical considerations to ensure “safe and effective” (P079 [30-39, Zambia, Public Health Data Analyst]) implementation. P079 [30-39, Zambia, Public Health Data Analyst] cautiously warned, “I think it’s important to be cautious about the use of AI and machine learning in healthcare. We need to make sure that we are using these technologies in a way that is ethical and that does not have negative consequences for patients”. Similarly P006 [30-39, Tanzania, Public Health Researcher] highlighted the importance of ethical data collection and analysis, stating “It is important to ensure that AI is used to collect and analyze data in a way that is ethical and respectful of African communities”.
@@
3.3.5 Theme: Capacity Building for Effective Implementation and Adoption of AI Solutions in Health in Africa
@@
Figure 11: Major theme from the expert IDIs: Capacity building. Summary of the subthemes illustrating the key capacity building initiatives to build an effective ecosystem for developing and adopting AI solutions in health in Africa .
@@
4 Discussion
Our study identified expert and general population perspectives on factors contributing to health inequities in Africa and the potential for AI applications in healthcare to improve population health and reduce health inequities in Africa. A major recurring theme identified relates to the need for contextually-grounded, culturally-aware, and community-driven approaches. For example, this includes the need to understand specific social and structural causes of health inequities, including those related to infrastructure, healthcare access, and local political, regulatory, and governmental environments Qoseem2024; Owoyemi2020; Oladipo2024; the need for a participatory approach to problem formulation grounded in the needs of the specific communities and contexts in which a system is intended to be used Martin2020; Suresh2024; Passi2019; as well as contextualized approaches to conceptualizing, measuring, and mitigating issues related to algorithmic fairness, bias, and transparency Sambasivan2021; Asiedu2024. Regarding the potential for AI to promote health equity, participants noted that AI systems and interventions designed to address specific clinical needs (e.g., tools for patient registration, clinical decision support, or diagnostics), structural barriers (e.g., healthcare access), and context-specificity (e.g., augmentation of LLM prompts) may be effective. Furthermore, participants highlighted that, in order to realize this potential, it is critical to address complex structural and social challenges related to infrastructure and capacity building, trust and community attitudes in the context of colonial history, as well as technical limitations related to biases in datasets and AI systems.
Limitations: A limitation of our work is the restriction of the population surveyed to one that possessed advanced English comprehension with a relatively high degree of technical literacy and education. We believed this to be necessary in order to engage most effectively with the survey. However, we acknowledge that this may limit the extent to which the findings of this study generalize.As such, it is important that future work extends our approach with more accessible methods such as on-the-ground verbal interactions that seek to understand the perspectives of those not well-represented in our study, and extend it to local languages. Furthermore, as our study primarily leverages qualitative insights from interviews and surveys, an important area of future research involves integration of these findings with real-world evidence of the relationship between specific axes of disparities and AI-enabled health applications.
5 Conclusion
Acknowledgements
The authors would like to acknowledge the insights and feedback from Maya Leventer-Roberts, Sunny Jansen, Chirag Nagpal and Vishwali Mhasawade. We would also like to thank Gutcheck for their rigor, timeliness and reach in administering and analyzing the general population surveys, and D’Well for their transcription services.
Data availability
The codebooks, and transcripts from in-depth interviews, cross tabs from surveys, and survey and in-depth interview questions are provided as part of this submission.
Author information
These authors contributed equally: Mercy Nyamewaa Asiedu, Iskandar Haykel and Awa Dieng
MA initiated and designed the overall study methodology and protocol, supervised and managed the progress of the study. KH and MA provided funding for the study. MA and IH developed the qualitative interview and quantitative study materials (IDI pre-surveys, interview guides, gen pop surveys, recruitment material, screening forms, etc). AD, SP, and NR provided feedback on the study material. IH, AD, and MA conducted the qualitative interviews - IH, AD administered the interviews, MA recruited and filtered participants for eligibility, and oversaw scheduling, incentives and transcription. KK designed the interview analysis methodology pipeline. IH, AD, KK, TA, FO, and CC analyzed the interview transcripts including thematic analysis, 2 rounds of axial coding, initial binning into sub themes and themes. AD finalized the interview codes and the themes, analyzed and interpreted the findings. MA finalized the survey analysis and interpreted the survey findings. MA, IH, AD, KK, TA, FO, NR, SP, and KH provided feedback throughout the analysis process. IH and FO developed the literature review. MA, IH, AD, KK, TA, and SP wrote the initial draft of the manuscript. All authors reviewed and provided feedback for the final draft of the manuscript.
Competing interests
This study was funded by Google LLC and/or subsidiary thereof (Google). All the authors are employees of Google and may own stock as a part of a standard compensation package.
[…]
{refsection}
Appendix
@@ Appendix A.1 Study materials
A.1.1 survey questions
A.1.2 In depth interview guides
Appendix A.2 Per country comparison of Gen Pop survey results
A.2.1 Per country comparison of Gen Pop survey results
Per country general population responses to likert scale questions on agreement with AI definitions, familiarity with AI, trust, benefits, concerns, country-level disparities, and colonialism by country. A/B/C/D/E Indicates statistical significance at 95% confidence
Figure A.1: Q. To what extent do you agree/disagree with the definition of AI? Q. You said that you [INSERT RESPONSE FROM PREVIOUS QUESTION] with the provided definition of AI (Artificial Intelligence). In your own words, how would you define AI (Artificial Intelligence)? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.2: Q. To what extent do you agree/disagree with the definition of Machine Learning? Q. You said that you [INSERT RESPONSE FROM PREVIOUS QUESTION] with the provided definition of Machine Learning. In your own words, how would you define Machine Learning? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.3: Q. How would you rate your familiarity with artificial intelligence (AI) and its associated use cases? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.4: Q. How would you describe your sense of certainty/uncertainty about the benefits of AI for societies and people in your country? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.5: Q. How would you describe your sense of certainty/uncertainty about the benefits of AI for societies and people in your country? Base: Total n=672
Figure A.6: Q. How would you rate your level of concern on the use of AI in health interventions in your country? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.7: Q. How likely are you to trust a health diagnosis for a health issue you have, if you learn that it was performed at least in part through the use of artificial intelligence? Q. How likely are you to trust a health intervention affecting you, e.g., a diagnosis for a health issue you have, if you learn that it was performed at least in part through the use of machine learning? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.8: Q. How likely are you to use a hypothetical AI tool called "Kasa" where you can provide your health information or ask health-related questions to Kasa and it would provide automated advice or responses about improving your health? Q. If you were told the Kasa tool performs less accurately for certain groups of people, would that change your answer that you are [INSERT RESPONSE FROM PREVIOUS QUESTION] to use the tool? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.9: Q. Do you see any connection between AI and colonialism? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.10: Q. How likely do you believe it is that AI will be a force for health empowerment in your country? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.11: Q. How likely do you believe it is that AI is likely to empower the High-Income Countries technologically and economically at the expense of countries in sub-Saharan Africa? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.12: Q. To what extent do you agree/disagree with the definition of fairness in AI? Q. You said that you [INSERT RESPONSE FROM PREVIOUS QUESTION] with the provided definition of fairness in AI (Artificial Intelligence). In your own words, how would you define fairness in AI (Artificial Intelligence)? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.13: Q. To what extent do you agree/disagree with the definition of bias in AI? Q. You said that you [INSERT RESPONSE FROM PREVIOUS QUESTION] with the provided definition of bias in AI (Artificial Intelligence). In your own words, how would you define bias in AI (Artificial Intelligence)? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.14: Q. Do you believe that an automated AI tool would be biased towards or against you? In other words, do you have any reason to believe that an AI-based tool would treat you better or worse than other people? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.15: Q. Would you be willing to participate in HealthMax’s project? Base: Total n=335, South Africa n=74, Kenya n=55, Nigeria n=76, Ghana n=61, Rwanda n=69 Q. Would you be willing to participate in HealthMax’s Afrique’s project? Base: Total n=337, South Africa n=54, Kenya n=70, Nigeria n=93, Ghana n=64, Rwanda n=56
Figure A.16: Q. Imagine you have a medical concern about your body, but it is difficult for you to easily get a medical check-up. You know of a widely used smartphone application that allows you to communicate via speech/text with a ChatBot trained in medical diagnosis and counseling. This ChatBot is capable not only of receiving information about medical concerns, but can also analyze and diagnose medical conditions from images and patients’ descriptions of symptoms. You have the following conversation with the chatbot.
Figure A.17: Q. To what extent do you believe that the answer to your question is in the response? Q. What are your thoughts on the length of information provided? Q. What are your thoughts on the accuracy of information provided? Q. Do you believe your diagnosis is in the list of the suggested conditions? Base: South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.18: Q. What do you do next? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.19: Q. Imagine you have a medical concern about your body, but it is difficult for you to easily get a medical check-up. You know of a widely used smartphone application that allows you to communicate via speech/text with a ChatBot trained in medical diagnosis and counseling. This ChatBot is capable not only of receiving information about medical concerns, but can also analyze and diagnose medical conditions from images and patients’ descriptions of symptoms. You have the following conversation with the chatbot.
Figure A.20: Q. To what extent do you believe the additional information provided by the AI/ML tool specific to country and location is helpful? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.21: Q. Which of the following recommended tasks would you do? Please click here to review the conversation Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.22: Q. Imagine you have a medical concern about your body, but it is difficult for you to easily get a medical check-up. You know of a widely used smartphone application that allows you to communicate via speech/text with a ChatBot trained in medical diagnosis and counseling. This ChatBot is capable not only of receiving information about medical concerns, but can also analyze and diagnose medical conditions from images and patients’ descriptions of symptoms. You have the following conversation with the chatbot.
Figure A.23: Q. To what extent do you believe the additional information on gender and religion you added to the question is helpful? Q. To what extent do you believe the additional response provided by the AI tool is helpful? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.24: Q. Which of the following recommended tasks would you do? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125

### v4

@@ A.1.3. Comparisons of expert and general population survey on AI axes of disparities
Experts were asked to rate how likely different potential axes of disparities would cause AI to perform differently for sub-groups in Africa while Gen Pop. respondents were asked to rate how likely different demographics would cause AI to perform differently for the participant themselves. Results are reported in Fig. A.3. Overall most experts (71%) indicated that all demographics were likely or very likely to be axes of AI disparities/AI biases for sub-groups in Africa, with literacy, race, language, rural vs. urban location, and national income level being the top 5 axes of disparities. Most Gen Pop. respondents did not feel that AI tools would perform differently for them compared to other people. In agreement with experts, Gen Pop. indicated that literacy had the highest level of causing an AI tool to perform differently. Given that our respondents were mostly from educated backgrounds, 32% indicated that their literacy level would allow an AI tool for health to perform better for them. Country of residence, country of birth, race, and colonial history of country of origin had the highest percentage of respondents indicating that they would cause an AI tool to perform worse for them (Fig. A.3).
@@
Figure A.3. Perceptions from a) experts and b) Gen Pop. respondents on how likely different demographics would cause an AI tool to perform differently for people in Africa. Note that Better is not necessarily good, it reflects that our participant population is better off than other people in the same country. e.g., Most of our survey participants are skewed towards high literacy and education levels compared to the general population, and so AI may perform better for them than for others.
Appendix A.2 Per country comparison of Gen Pop survey results
A.2.1. Per country comparison of Gen Pop survey results
Per country general population responses to likert scale questions on agreement with AI definitions, familiarity with AI, trust, benefits, concerns, country-level disparities, and colonialism by country. A/B/C/D/E Indicates statistical significance at 95% confidence.
Figure A.4. Q. To what extent do you agree/disagree with the definition of AI? Q. You said that you [INSERT RESPONSE FROM PREVIOUS QUESTION] with the provided definition of AI (Artificial Intelligence). In your own words, how would you define AI (Artificial Intelligence)? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.5. Q. To what extent do you agree/disagree with the definition of Machine Learning? Q. You said that you [INSERT RESPONSE FROM PREVIOUS QUESTION] with the provided definition of Machine Learning. In your own words, how would you define Machine Learning? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.6. Q. How would you rate your familiarity with artificial intelligence (AI) and its associated use cases? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.7. Q. How would you describe your sense of certainty/uncertainty about the benefits of AI for societies and people in your country? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.8. Q. How would you describe your sense of certainty/uncertainty about the benefits of AI for societies and people in your country? Base: Total n=672
Figure A.9. Q. How would you rate your level of concern on the use of AI in health interventions in your country? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.10. Q. How likely are you to trust a health diagnosis for a health issue you have, if you learn that it was performed at least in part through the use of artificial intelligence? Q. How likely are you to trust a health intervention affecting you, e.g., a diagnosis for a health issue you have, if you learn that it was performed at least in part through the use of machine learning? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.11. Q. How likely are you to use a hypothetical AI tool called ”Kasa” where you can provide your health information or ask health-related questions to Kasa and it would provide automated advice or responses about improving your health? Q. If you were told the Kasa tool performs less accurately for certain groups of people, would that change your answer that you are [INSERT RESPONSE FROM PREVIOUS QUESTION] to use the tool? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.12. Q. Do you see any connection between AI and colonialism? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.13. Q. How likely do you believe it is that AI will be a force for health empowerment in your country? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.14. Q. How likely do you believe it is that AI is likely to empower the High-Income Countries technologically and economically at the expense of countries in sub-Saharan Africa? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.15. Q. To what extent do you agree/disagree with the definition of fairness in AI? Q. You said that you [INSERT RESPONSE FROM PREVIOUS QUESTION] with the provided definition of fairness in AI (Artificial Intelligence). In your own words, how would you define fairness in AI (Artificial Intelligence)? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.16. Q. To what extent do you agree/disagree with the definition of bias in AI? Q. You said that you [INSERT RESPONSE FROM PREVIOUS QUESTION] with the provided definition of bias in AI (Artificial Intelligence). In your own words, how would you define bias in AI (Artificial Intelligence)? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.17. Q. Do you believe that an automated AI tool would be biased towards or against you? In other words, do you have any reason to believe that an AI-based tool would treat you better or worse than other people? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.18. Q. Would you be willing to participate in HealthMax’s project? Base: Total n=335, South Africa n=74, Kenya n=55, Nigeria n=76, Ghana n=61, Rwanda n=69 Q. Would you be willing to participate in HealthMax’s Afrique’s project? Base: Total n=337, South Africa n=54, Kenya n=70, Nigeria n=93, Ghana n=64, Rwanda n=56
Figure A.19. Q. Imagine you have a medical concern about your body, but it is difficult for you to easily get a medical check-up. You know of a widely used smartphone application that allows you to communicate via speech/text with a ChatBot trained in medical diagnosis and counseling. This ChatBot is capable not only of receiving information about medical concerns, but can also analyze and diagnose medical conditions from images and patients’ descriptions of symptoms. You have the following conversation with the chatbot.
Figure A.20. Q. To what extent do you believe that the answer to your question is in the response? Q. What are your thoughts on the length of information provided? Q. What are your thoughts on the accuracy of information provided? Q. Do you believe your diagnosis is in the list of the suggested conditions? Base: South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.21. Q. What do you do next? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.22. Q. Imagine you have a medical concern about your body, but it is difficult for you to easily get a medical check-up. You know of a widely used smartphone application that allows you to communicate via speech/text with a ChatBot trained in medical diagnosis and counseling. This ChatBot is capable not only of receiving information about medical concerns, but can also analyze and diagnose medical conditions from images and patients’ descriptions of symptoms. You have the following conversation with the chatbot.
Figure A.23. Q. To what extent do you believe the additional information provided by the AI/ML tool specific to country and location is helpful? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.24. Q. Which of the following recommended tasks would you do? Please click here to review the conversation Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.25. Q. Imagine you have a medical concern about your body, but it is difficult for you to easily get a medical check-up. You know of a widely used smartphone application that allows you to communicate via speech/text with a ChatBot trained in medical diagnosis and counseling. This ChatBot is capable not only of receiving information about medical concerns, but can also analyze and diagnose medical conditions from images and patients’ descriptions of symptoms. You have the following conversation with the chatbot.
Figure A.26. Q. To what extent do you believe the additional information on gender and religion you added to the question is helpful? Q. To what extent do you believe the additional response provided by the AI tool is helpful? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
Figure A.27. Q. Which of the following recommended tasks would you do? Base: Total n=672, South Africa n=128, Kenya n=125, Nigeria n=169, Ghana n=125, Rwanda n=125
A.2.2. Thematic analysis of expert qualitative interviews
A.2.2.1. Theme: Trust In AI in Africa
Table A.1. A summary of the theme and subthemes from the expert in-depth interviews
@@ A.2.2.1. Theme: Trust In AI in Africa
Figure A.28. Major theme from the expert IDIs: Trust in AI. Summary of the subthemes illustrating the identified causes for mistrust and approaches for building trust.
Subtheme: Mistrust due to marginalization, colonization, misunderstanding of the technology, and perceived lack of transparency, fairness, and accountability Mistrust due to marginalization: Experts pointed to skepticism surrounding AI in Africa based on a history of marginalization from the Global North. “a lot of times things that are seen to be coming from the Global North are viewed with a lot of skepticism. -P025 [50-59, Nigeria, Medical Doctor]. This distrust is further fueled by concerns of exploitation by foreign entities and worries of job insecurity among healthcare workers with sentiments, such as “What are the hidden things that they really want that they’re not letting us see?”- P025 [50-59, Nigeria, Medical Doctor], or “people are scared of the power of AI, they think that it is going to wipe their businesses off, is going to take their jobs” - P168 [30-39, Ghana, Educator].
@@
A.2.2.2. Theme: Opportunities for AI to Address Health in Africa
@@
Table A.2. A summary of the theme and subthemes from the expert in-depth interviews
THEME: Opportunities for AI to Address Health in Africa
Subtheme Representative quote
Figure A.29. Major theme from the expert IDIs: Opportunities for AI to address health in Africa. Summary of the subthemes illustrating the foremost use cases for AI in health and the desired features for AI solutions to improve efficacy and adoption.
@@
A.2.2.3. Theme: Intersectional Inequities and Systemic Factors Impacting Health Equity in Africa
@@
Table A.3. A summary of the theme and subthemes from the expert in-depth interviews
THEME: Intersectional Inequities and Systemic Factors
Impacting Health Equity in Africa
Subtheme Representative quote
Figure A.30. Major theme from the expert IDIs: Intersectional Inequities and Systemic Factors impacting health equity in Africa. Summary of the subthemes illustrating the existing causes of health inequities and proposed approaches for addressing them.
@@
A.2.2.4. Theme: Social and Ethical Considerations for Developing and Deploying AI Solutions in Health in Africa
@@
Table A.4. A summary of the theme and subthemes with representative quotes from the expert in-depth interviews
THEME: Social and Ethical Considerations for Developing
and Deploying AI solutions in Health in Africa
Subtheme Representative quote
Figure A.31. Major theme from the expert IDIs: Social and Ethical considerations for developing and deploying AI solutions in Health in Africa. Summary of the subthemes illustrating the social, ethical, privacy and data-related considerations.
@@
In addition to social considerations, participants emphasized the importance of responsible and ethical implementation of AI solutions to ensure fairness and avoid negative consequences. Ethical deployment: While the participants acknowledged that AI holds great promise for revolutionizing healthcare in Africa, potentially making it more “accessible, efficient, and equitable” (P115 [21-29, Cameroon, Machine Learning Researcher]), they caution against uncritical adoption, stressing the need for ethical considerations to ensure “safe and effective” (P079 [30-39, Zambia, Public Health Data Analyst]) implementation. P079 [30-39, Zambia, Public Health Data Analyst] cautiously warned, “I think it’s important to be cautious about the use of AI and machine learning in healthcare. We need to make sure that we are using these technologies in a way that is ethical and that does not have negative consequences for patients”. Similarly P006 [30-39, Tanzania, Public Health Researcher] highlighted the importance of ethical data collection and analysis, stating “It is important to ensure that AI is used to collect and analyze data in a way that is ethical and respectful of African communities”.
@@
A.2.2.5. Theme: Capacity Building for Effective Implementation and Adoption of AI Solutions in Health in Africa
@@
Table A.5. A summary of the theme and subthemes with representative quotes from the expert in-depth interviews
THEME: Capacity Building for Effective Implementation and Adoption
of AI Solutions in Health in Africa
Subtheme Representative quote
Figure A.32. Major theme from the expert IDIs: Capacity building. Summary of the subthemes illustrating the key capacity building initiatives to build an effective ecosystem for developing and adopting AI solutions in health in Africa .
@@ Appendix A.3 Study materials
A.3.1. survey questions
A.3.2. In depth interview guides